USE METHOD OF SUBSTITUTION TO SOLVE SYSTEM OF LINEAR EQUATION
7X+5Y=7/6
Y=6X-1

Answer by lawdej(58)

7X+5Y=7/6
Y=6X-1
substitute the second one to the first equation
7x+5(6x-1)=7/6
7x+30x-5=7/6
37x=7/6+5
37x=37/6
x=1/6
to find the value of y substitute x=1/6 to y=6x-1
y= 6(1/6)-1
y=1-1
y=0